# Policy-Driven World Model Adaptation for Robust Offline Model-based Reinforcement Learning

### v2

@@ 6 Experimental Results
In this section, we benchmark our algorithm (ROMBRL) against a range of state-of-the-art (SOTA) offline RL baselines across two task sets comprising a total of 15 continuous control tasks. The first task set is the widely used D4RL MuJoCo suite [DBLP:journals/corr/abs-2004-07219], which includes three types of robotic agents, each with offline datasets of four different quality levels. Since D4RL MuJoCo features deterministic dynamics, we further evaluate on a more challenging set of Tokamak control tasks, where the underlying dynamics are highly stochastic. To assess the robustness of the policies learned by different algorithms, we introduce noise to the environments’ dynamics during deployment. Specifically, at each time step $t$, the state change (in each dimension) $\Delta s_{t}=s_{t+1}-s_{t}$ is perturbed with zero-mean Gaussian noise, where the standard deviation is proportional to $\Delta s_{t}$. In our experiments, we use 5% measurement noise.
@@
Table 1: Comparison of our algorithm with state-of-the-art offline RL methods on the D4RL MuJoCo benchmark. The abbreviations ‘hc’, ‘hp’, and ‘wk’ denote HalfCheetah, Hopper, and Walker2d, respectively. To evaluate robustness, we introduce measurement noise into the real MuJoCo dynamics. Each value represents the normalized score, as proposed in [DBLP:journals/corr/abs-2004-07219], of the policy trained by the corresponding algorithm. These scores are undiscounted returns normalized to approximately range between 0 and 100, where a score of 0 corresponds to a random policy and a score of 100 corresponds to an expert-level policy. For each algorithm, we report the average score of the final 100 policy learning epochs and its standard deviation across three random seeds. The best and second-best results for each task are bolded and marked with * respectively. The last row includes Cohen’s $d$ to indicate the significance of our algorithm’s improvement over other methods.
@@
Data Type Agent Type ROMBRL (ours) CQL EDAC COMBO RAMBO MOBILE
random hc 39.3* (4.0) 18.3 (1.2) 14.9 (7.8) 5.8 (2.1) 36.7 (3.4) 40.3 (2.1)
random hp 31.3 (0.1) 10.1 (0.4) 14.1 (5.6) 10.3 (3.4) 25.7* (6.1) 22.7 (9.1)
random wk 21.7 (0.1) 2.3 (1.9) 0.7 (0.3) 0.0 (0.1) 13.1 (3.0) 17.2* (3.3)
medium hc 77.5* (1.7) 48.5 (0.1) 62.1 (1.0) 49.6 (0.2) 78.1 (1.2) 72.8 (0.5)
medium hp 102.6 (2.6) 56.1 (0.4) 54.8 (38.4) 71.1 (3.1) 71.7 (8.2) 89.9* (16.1)
medium wk 79.4* (4.2) 73.9 (0.4) 83.6 (1.4) 72.9 (3.1) 13.7 (5.0) 71.5 (6.1)
med-rep hc 73.8 (0.4) 45.7 (0.3) 53.2 (0.7) 46.4 (0.2) 65.8 (0.4) 67.6* (4.2)
med-rep hp 105.3 (0.7) 94.1 (4.7) 101.3 (0.6) 99.5 (1.3) 97.6 (1.1) 104.9* (1.8)
med-rep wk 90.5 (2.4) 72.2 (17.3) 83.1 (0.3) 71.7 (1.9) 80.9 (1.0) 85.2* (2.3)
med-exp hc 88.9* (2.1) 88.3 (0.5) 61.2 (6.9) 90.5 (0.0) 83.4 (1.4) 82.9 (5.2)
med-exp hp 111.9 (0.1) 109.1 (0.9) 55.2 (41.3) 110.1* (0.5) 83.5 (4.9) 79.6 (45.5)
med-exp wk 110.7* (2.6) 109.9 (0.1) 60.6 (46.3) 37.8 (50.9) 19.5 (20.7) 113.7 (2.5)
Average Score 77.7 (0.5) 60.7 (1.2) 53.7 (4.6) 55.5 (3.6) 55.8 (1.3) 70.7* (2.4)
Cohen’s $d$ - 18.9 7.4 8.5 22.9 4.1
@@ 6.1 Results on D4RL MuJoCo
Here, we compare ROMBRL with several representative offline RL methods, including two model-free algorithms: CQL [DBLP:conf/nips/KumarZTL20], EDAC [DBLP:conf/nips/AnMKS21], and three model-based algorithms: COMBO [DBLP:conf/nips/YuKRRLF21], RAMBO [DBLP:conf/nips/RigterLH22], MOBILE [DBLP:conf/icml/SunZJLYY23]. These baselines are carefully selected: EDAC and MOBILE are top model-free and model-based offline RL algorithms on the D4RL Mujoco benchmark, according to [DBLP:conf/icml/SunZJLYY23]; CQL and its model-based extension, COMBO, are widely adopted in real-life robotic tasks; RAMBO is explicitly designed to enhance robustness in offline RL.
In Table 1, we report the convergent performance of each algorithm on all tasks. Following the protocol in the D4RL benchmarking paper [DBLP:journals/corr/abs-2004-07219], convergent performance is defined as the mean evaluation score over the final 100 policy learning epochs. We present the mean (standard deviation) across runs with different random seeds. ROMBRL ranks first on 7 out of 12 tasks and second on the remaining 5. In terms of average performance, ROMBRL significantly outperforms all baselines. To quantify the improvement in average performance, we compute Cohen’s $d$ between ROMBRL and each baseline. According to Cohen’s rule of thumb [sawilowsky2009new], a value of $d\geq 2$ indicates a very large and statistically significant difference. These results demonstrate the superior performance of our algorithm in robust deployment. Additionally, in Figure 2, we present the training curves of each algorithm on all D4RL MuJoCo tasks.
@@ 6.2 Results on Tokamak Control
We further evaluate our algorithm on three target tracking tasks for tokamak control. The tokamak is one of the most promising confinement devices for achieving controllable nuclear fusion, where the primary challenge lies in confining the plasma, i.e., an ionized gas of hydrogen isotopes, while heating it and increasing its pressure to initiate and sustain fusion reactions [1512794]. Tokamak control involves applying a series of direct actuators (e.g., ECH power, magnetic field) and indirect actuators (e.g., setting targets for the plasma shape and density) to confine the plasma to achieve a desired state or track a given target. This sophisticated physical process is an ideal test bed for our algorithm.
As shown in Figure 3, we use a well-trained data-driven dynamics model provided by [DBLP:journals/corr/abs-2404-12416] as a "ground truth" simulator for the nuclear fusion process during evaluation, and generate a dataset containing 111305 transitions for offline RL. We select reference shots (each of which represents a fusion process) from DIII-D [^6] DIII-D is a tokamak device located in San Diego, California, operated by General Atomics., and use trajectories of Ion Rotation, Electron Density, and $\beta_{N}$ within them as targets for three tracking tasks. These are critical quantities in tokamak control, particularly $\beta_{N}$, which serves as an economic indicator of the efficiency of nuclear fusion. These continuous control tasks are highly stochastic, as the underlying dynamics model is an ensemble of recurrent probabilistic neural networks (RPNNs) and each state transition is a sample from this model. For details about the simulator, and the design of the state/action spaces and reward functions, please refer to Appendix K.
Table 2: Comparison of our algorithm with state-of-the-art offline RL methods on the Tokamak Control benchmark. To evaluate robustness, we inject measurement noise into the tokamak dynamics. Each value represents the negative episodic tracking error of a specific physical quantity in the reference shot. For each algorithm, we report the average evaluation performance of the final 100 policy learning epochs and its standard deviation across three random seeds. The best and second-best results for each task are bolded and marked with * respectively. The last row reports Cohen’s $d$ to quantify the significance of our algorithm’s improvement over competing methods. A value of $d\geq 2$ indicates a statistically significant improvement.
@@
In addition to the baselines presented in Table 1, we include a recent model-based offline RL method, BAMCTS [DBLP:journals/corr/abs-2410-11234], which has also been evaluated on Tokamak Control tasks. The benchmarking results are shown in Table 2. The evaluation metric is the negative episodic tracking error of the reference shots, computed as the sum of mean squared errors between the achieved and target quantities at each time step. As in previous experiments, we report the average score over the final 100 policy learning epochs as the policy’s convergent performance. The results show that ROMBRL ranks first in 2 out of 3 tasks and second in the remaining one. Moreover, ROMBRL achieves the best average performance across all three tasks, with notably low variance. Cohen’s $d$ further confirms that ROMBRL’s improvement over each baseline is statistically significant. In contrast, the performance of several SOTA model-based methods: RAMBO, MOBILE, and BAMCTS, drops significantly in this stochastic and noisy benchmark, showing poor robustness and high variance across runs. The training curves of each algorithm for all tasks are shown as Figure 4 in Appendix K.
@@ 7 Conclusion
Acknowledgments
This work was funded in part by Department of Energy Fusion Energy Sciences under grant DE-SC0024544.
@@ Appendix A A Discussion on $\omega_{1}^{k}$ in Equation (4)
Appendix B Proof of Theorem 1
We begin our proof by presenting key lemmas. The first one is Lemma 25 from [DBLP:conf/nips/AgarwalKKS20].
Lemma 1.
@@ Appendix B Proof of Theorem 1
The second lemma is analogous to the simulation lemma [DBLP:conf/icml/RossB12] but does not assume knowledge of the reward function.
Lemma 2.
@@
Now, we formally begin our proof of Theorem 1:
@@
Starting from Lemma 2, we have: ($f_{\phi}(s,a)=TV(P_{\phi*}(\cdot|s,a),P_{\phi}(\cdot|s,a))^{2}$)
@@
According to Lemma 1, we further have:
@@
Appendix C Proof of Theorem 2
For tabular MDPs, the maximum likelihood estimator $P_{\bar{\phi}}(\cdot|s,a)$ is given by the empirical distribution derived from the samples in $\mathcal{D}_{\mu}$ at each state-action pair $(s,a)$, i.e., $P_{\hat{\phi}}(\cdot|s,a)$. Thus, we have: $\mathbb{E}_{(s,a)\sim\mathcal{D}_{\mu},(r,s^{\prime})\sim P_{\bar{\phi}}(\cdot|s,a)}\left[\log P_{\bar{\phi}}(r,s^{\prime}|s,a)-\log P_{\phi^{*}}(r,s^{\prime}|s,a)\right]=\mathbb{E}_{(s,a)\sim\mathcal{D}_{\mu}}\left[KL(P_{\hat{\phi}}(\cdot|s,a)||P_{\phi^{*}}(\cdot|s,a))\right]$. According to [mardia2020concentration], when $3\leq K\leq\frac{N_{sa}C_{0}}{e}+2$:
@@ Appendix C Proof of Theorem 2
Appendix D Proof of Theorem 3
We begin by listing the key lemmas used in the proof. The first lemma is from [laurent2000adaptive], which claims:
Lemma 3.
@@ Appendix D Proof of Theorem 3
Let $X_{1}=\frac{(\mu-\hat{\mu})^{2}}{\sigma^{2}/n}$, then $X_{1}\sim\chi^{2}_{1}$. According to Lemma 3, $\mathbb{P}(X\geq n+2\sqrt{nt}+2t)\leq\exp(-t),\ \forall t>0$. Let $\exp(-t)=\frac{\delta^{\prime}}{2}$, we obtain:
@@
Further, we observe that $\frac{\hat{\sigma}^{2}}{\sigma^{2}}=\frac{\sum_{i=1}^{n}(x_{i}-\hat{\mu})^{2}}{n\sigma^{2}}=\frac{X_{2}}{n}$, where $X_{2}\sim\chi^{2}_{n-1}$. Applying the two formulas in Lemma 3 and letting $\exp(-t)=\delta^{\prime}/4$, we have:
@@
Since $P_{\bar{\phi}}$ is an MLE of $P_{\phi^{*}}$, $\bar{\mu}_{sa}^{i}=\hat{\mu}_{sa}^{i}$ and $\bar{\sigma}_{sa}^{2,i}=\hat{\sigma}_{sa}^{2,i}$ are the sample mean and sample variance, respectively. Applying Lemma 4 and letting $\delta^{\prime}=\frac{\delta}{2\mathcal{D}_{sa}^{\mu}(d+1)}$, we obtain:

### v3

@@ 6 Experimental Results
Table 1: Comparison with SOTA offline RL methods on the D4RL MuJoCo benchmark. The abbreviations ‘hc’, ‘hp’, and ‘wk’ denote HalfCheetah, Hopper, and Walker2d, respectively. To evaluate robustness, we add measurement noise to the real MuJoCo dynamics. Each value represents the normalized score, as proposed in Fu et al. (2020), of the policy trained by the corresponding algorithm. These scores are undiscounted returns normalized to approximately range between 0 and 100, where a score of 0 corresponds to a random policy and a score of 100 corresponds to an expert-level policy. For each algorithm, we report the average score of the final 100 policy learning epochs and its standard deviation across three random seeds. The best and second-best results for each task are bolded and marked with * respectively. The last row includes Cohen’s $d$ to indicate the significance of our algorithm’s improvement over other methods.
@@
Data Type Agent Type ROMBRL (ours) CQL EDAC COMBO RAMBO MOBILE RORL TRACER RFQI
random hc 39.3* (4.0) 18.3 (1.2) 14.9 (7.8) 5.8 (2.1) 36.7 (3.4) 40.3 (2.1) 29.0 (1.4) 10.8 (3.9) 2.8 (0.2)
random hp 31.3 (0.1) 10.1 (0.4) 14.1 (5.6) 10.3 (3.4) 25.7* (6.1) 22.7 (9.1) 10.2 (4.1) 5.7 (2.9) 3.7 (1.5)
random wk 21.7 (0.1) 2.3 (1.9) 0.7 (0.3) 0.0 (0.1) 13.1 (3.0) 17.2* (3.3) 0.0 (0.2) 8.9 (3.6) 7.0 (0.6)
medium hc 77.5* (1.7) 48.5 (0.1) 62.1 (1.0) 49.6 (0.2) 78.1 (1.2) 72.8 (0.5) 60.3 (1.3) 47.7 (0.6) 40.3 (0.3)
medium hp 102.6 (2.6) 56.1 (0.4) 54.8 (38.4) 71.1 (3.1) 71.7 (8.2) 89.9 (16.1) 100.1* (8.5) 62.1 (5.4) 41.0 (4.9)
medium wk 79.4 (4.2) 73.9 (0.4) 83.6* (1.4) 72.9 (3.1) 13.7 (5.0) 71.5 (6.1) 89.9 (4.8) 50.9 (6.3) 55.4 (1.9)
med-rep hc 73.8 (0.4) 45.7 (0.3) 53.2 (0.7) 46.4 (0.2) 65.8 (0.4) 67.6* (4.2) 51.4 (1.5) 39.0 (2.3) 31.6 (2.6)
med-rep hp 105.3 (0.7) 94.1 (4.7) 101.3 (0.6) 99.5 (1.3) 97.6 (1.1) 104.9* (1.8) 102.0 (1.0) 63.6 (7.7) 14.6 (11.5)
med-rep wk 90.5 (2.4) 72.2 (17.3) 83.1 (0.3) 71.7 (1.9) 80.9 (1.0) 85.2 (2.3) 85.3* (2.1) 8.3 (5.5) 16.0 (4.5)
med-exp hc 88.9* (2.1) 88.3 (0.5) 61.2 (6.9) 90.5 (0.0) 83.4 (1.4) 82.9 (5.2) 85.2 (2.4) 73.3 (5.0) 43.5 (1.9)
med-exp hp 111.9 (0.1) 109.1 (0.9) 55.2 (41.3) 110.1* (0.5) 83.5 (4.9) 79.6 (45.5) 24.9 (0.1) 82.9 (43.7) 46.6 (1.8)
med-exp wk 110.7* (2.6) 109.9 (0.1) 60.6 (46.3) 37.8 (50.9) 19.5 (20.7) 113.7 (2.5) 109.2 (9.1) 76.0 (37.6) 68.6 (5.0)
Average Score 77.7 (0.5) 60.7 (1.2) 53.7 (4.6) 55.5 (3.6) 55.8 (1.3) 70.7* (2.4) 62.3 (0.2) 44.1 (4.8) 30.9 (2.1)
Cohen’s $d$ - 18.9 7.4 8.5 22.9 4.1 40.4 9.8 30.7
We benchmark our algorithm (ROMBRL) against a range of SOTA offline RL baselines across two task sets comprising 15 continuous control tasks. First, the widely used D4RL MuJoCo suite (Fu et al., 2020), covering three types of robotic agents, each with offline datasets of four different quality levels. Given D4RL’s deterministic dynamics, we further evaluate on the challenging, highly stochastic Tokamak control tasks. To assess the robustness of the policies learned by different algorithms, we add noise to the environments’ dynamics during deployment. Specifically, at each time step $t$, the state change (in each dimension) $\Delta s_{t}=s_{t+1}-s_{t}$ is perturbed with zero-mean Gaussian noise, where the standard deviation is proportional to $\Delta s_{t}$. In our experiments, we use 5% measurement noise.
For complete details on our experimental setup, including random seed selection and hyperparameter settings, please refer to Appendix M. For additional results, including computational cost analysis, performance under varying noise levels and training curves, see Appendix N.
@@ 6.1 Results on D4RL MuJoCo
Here, we compare ROMBRL with a comprehensive suite of baselines. First, we include several representative standard offline RL methods: two model-free algorithms, CQL (Kumar et al., 2020) and EDAC (An et al., 2021), and two model-based algorithms, COMBO (Yu et al., 2021) and MOBILE (Sun et al., 2023). These baselines are carefully selected: EDAC and MOBILE are recognized as top-performing model-free and model-based algorithms on the D4RL MuJoCo benchmark, respectively, according to Sun et al. (2023); meanwhile, CQL and its model-based extension, COMBO, are widely adopted in real-life robotic tasks.
Second, to rigorously evaluate robustness, we compare against four specialized robust offline RL algorithms: RAMBO (Rigter et al., 2022), RORL (Yang et al., 2022a), RFQI (Panaganti et al., 2022), and TRACER (Yang et al., 2024). The inclusion of these baselines is vital for a comprehensive assessment: RAMBO is explicitly designed to enhance robustness in offline MBRL via adversarial dynamics; Tracer, RORL, RFQI represent state-of-the-art robust model-free Offline RL approaches. Comparing against this diverse set allows us to demonstrate ROMBRL’s superiority over methods that rely on passive smoothing or value-based min-max optimization.
Table 2: Comparison of our algorithm against SOTA offline RL methods on standard (noiseless) and noisy D4RL MuJoCo benchmarks. To ensure a fair and reproducible comparison, baseline results are sourced from the official OfflineRL-Kit repository, which reports scores on 9 specific tasks. Consequently, all mean scores presented here are averaged over these 9 tasks, which are composed of 3 environments (HalfCheetah, Hopper, Walker2d) across 3 dataset types (medium, medium-replay, and medium-expert). The ’Performance Drop’ metric highlights the superior robustness of our method under deployment noise.
Metric ROMBRL (ours) CQL EDAC COMBO RAMBO MOBILE RORL
Mean Score (Standard Env.) 92.8 80.4 93.0 89.3 82.7 95.9 89.5
Mean Score (Noisy Env.) 93.4 77.5 68.3 72.2 66.0 85.3 78.7
Performance Drop (%) $\downarrow$ -0.6% 3.6% 26.6% 19.1% 20.2% 11.1% 12.1%
In Table 1, we report the convergent performance of each algorithm on all tasks. Following the protocol in the D4RL benchmarking paper (Fu et al., 2020), convergent performance is defined as the mean evaluation score over the final 100 policy learning epochs. We present the mean (standard deviation) across runs with different random seeds. ROMBRL ranks first on 7 out of 12 tasks and second on 4 tasks. In terms of average performance, ROMBRL significantly outperforms all baselines. To quantify the improvement in average performance, we compute Cohen’s $d$ between ROMBRL and each baseline. According to Cohen’s rule of thumb (Sawilowsky, 2009), a value of $d\geq 2$ indicates a very large and statistically significant difference. These results demonstrate the superior performance of our algorithm in robust deployment.
Table 2 quantitatively demonstrates the core trade-off and ultimate benefit of our robust design. While ROMBRL’s performance on the standard (noiseless) benchmark is highly competitive, on par with the SOTA method MOBILE and significantly outperforming the robust baseline RAMBO, it uniquely achieves the top score under noisy deployment conditions where other baselines falter. This slight performance difference in the ideal, noiseless setting is an expected and theoretically justified consequence of our learning objective (Eq. 5), which explicitly optimizes for worst-case robustness across an uncertainty set of models rather than overfitting to a single, known environment. Ultimately, these results validate that ROMBRL provides a powerful solution to the sim-to-real challenge by drastically improving deployment robustness—as shown by its minimal performance drop—without a significant compromise in standard benchmark performance.
@@ 6.2 Results on Tokamak Control
We further evaluate our algorithm on three target tracking tasks for tokamak control. The tokamak is one of the most promising confinement devices for achieving controllable nuclear fusion, where the primary challengec is confining the plasma, i.e., an ionized gas of hydrogen isotopes, while heating it and increasing its pressure to initiate and sustain fusion reactions (Pironti and Walker, 2005). Tokamak control involves applying a direct actuators (e.g., ECH power, magnetic field) and indirect actuators (e.g., setting targets for the plasma shape and density) to confine the plasma to achieve a desired state or track a given target. This sophisticated physical process is an ideal test bed for our algorithm.
As shown in Figure 2, we use a well-trained data-driven dynamics model provided by Char et al. (2024) as a "ground truth" simulator for the nuclear fusion process during evaluation, and generate a dataset containing 111305 transitions for offline RL. We select reference shots (each of which represents a fusion process) from DIII-D [^6] DIII-D is a tokamak device located in San Diego, California, operated by General Atomics., and use trajectories of Ion Rotation, Electron Density, and $\beta_{N}$ within them as targets for three tracking tasks. These are critical quantities in tokamak control, particularly $\beta_{N}$, which serves as an economic indicator of the efficiency of nuclear fusion. These continuous control tasks are highly stochastic, as the underlying dynamics model is an ensemble of recurrent probabilistic neural networks (RPNNs) and each state transition is a sample from this model. For details about the simulator, and the design of the state/action spaces and reward functions, please refer to Appendix K.
Table 3: Comparison with SOTA offline RL methods on the Tokamak Control benchmark. We inject measurement noise to evaluate robustness. Each value represents the negative episodic tracking error of a specific physical quantity in the reference shot. For each algorithm, we report the average evaluation performance of the final 100 policy learning epochs and its standard deviation across three random seeds. The best and second-best results for each task are bolded and marked with * respectively. Last row shows Cohen’s $d$ quantifying improvement significance. A value of $d\geq 2$ indicates a statistically significant improvement.
@@
In addition to the baselines presented in Table 1, we include a recent model-based offline RL method, BAMCTS (Chen et al., 2024a), which has also been evaluated on Tokamak Control tasks. The benchmarking results are shown in Table 3. The evaluation metric is the negative episodic tracking error of the reference shots, computed as the sum of mean squared errors between the achieved and target quantities at each time step. As in previous experiments, we report the average score over the final 100 policy learning epochs as the policy’s convergent performance. The results show that ROMBRL ranks first in 2 out of 3 tasks and second in the remaining one. Moreover, ROMBRL achieves the best average performance across all three tasks, with notably low variance. Cohen’s $d$ further confirms that ROMBRL’s improvement over each baseline is statistically significant. In contrast, the performance of several SOTA model-based methods: RAMBO, MOBILE, and BAMCTS, drops significantly in this stochastic and noisy benchmark, showing poor robustness and high variance across runs. The training curves of each algorithm for all tasks are shown as Figure 7 in Appendix K.
6.3 Ablation Study
In this section, we investigate the contribution of the proposed Stackelberg learning dynamics and the sensitivity of the algorithm to hyperparameter choices.
Effectiveness of Constrained Stackelberg Gradients.
A core theoretical contribution of ROMBRL is the derivation of the update rule in Eq. (9), which solves the constrained Stackelberg game. To validate its necessity, we designed an ablation study comparing three gradient update mechanisms: (1)Naive Alternating (Eq. (7)): Updates $\pi_{\theta}$ and $P_{\phi}$ alternately without lookahead, similar to the strategy used in RAMBO (Rigter et al., 2022). (2)Unconstrained Stackelberg (Eq. (8)): Applies implicit differentiation to the relaxed objective, accounting for model adaptation but ignoring the dynamics of the constraint enforcement (i.e., treating $\lambda$ as fixed). (3)Constrained Stackelberg (Ours, Eq. (9)): Fully accounts for the curvature of the constraint surface via the primal-dual dynamics of $\lambda$.
As illustrated in Figure 3, the results are compelling. The Naive approach yields the lowest return. Crucially, simply applying Stackelberg dynamics to the unconstrained objective (Eq. (8)) provides only a marginal improvement. However, our full method (Eq. (9)) achieves a substantial performance leap. This empirical evidence highlights a critical insight: robustness in constrained RL stems not merely from anticipating the model parameters ($\phi$) adaptation, but from anticipating the boundary adjustments of the uncertainty set (governed by $\lambda$). Only by incorporating the second-order information of the constraint (via matrix $H$ in Eq. (9)) can the policy effectively navigate the trade-off between reward maximization and robust conservatism.
Figure 3: Comparison of average returns on walker2d-medium using different update rules. The proposed Constrained Stackelberg update (Eq. (9)) significantly outperforms both Naive Alternating (Eq. (7)) and Unconstrained Stackelberg (Eq. (8)) approaches, verifying the necessity of accounting for constraint dynamics.
Hyperparameter Sensitivity.
We further examined the robustness of ROMBRL regarding the uncertainty radius $\epsilon$ and the learning rates. The results indicate that the choice of $\epsilon$ should correlate with the magnitude of the environmental disturbance, and the policy learning rate $\eta_{\theta}$ should be less than model learning rate $\eta_{\phi}$. Due to space constraints, the detailed quantitative results are provided in Appendix L.
@@ 7 Conclusion
Impact Statement
This paper presents work whose goal is to advance the field of machine learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ Appendix A A Discussion on $\omega_{1}^{k}$ in Equation (4)
Appendix B Proof of Theorem 3.1
We begin our proof by presenting key lemmas. The first one is Lemma 25 from (Agarwal et al., 2020).
Lemma B.1.
@@
The second lemma is analogous to the simulation lemma (Ross and Bagnell, 2012) but does not assume knowledge of the reward function.
Lemma B.2.
@@
Now, we formally begin our proof of Theorem 3.1:
@@
Starting from Lemma B.2, we have: ($f_{\phi}(s,a)=TV(P_{\phi*}(\cdot|s,a),P_{\phi}(\cdot|s,a))^{2}$)
@@
According to Lemma B.1, we further have:
@@
Appendix C Proof of Theorem 3.2
For tabular MDPs, the maximum likelihood estimator $P_{\bar{\phi}}(\cdot|s,a)$ is given by the empirical distribution derived from the samples in $\mathcal{D}_{\mu}$ at each state-action pair $(s,a)$, i.e., $P_{\hat{\phi}}(\cdot|s,a)$. Thus, we have: $\mathbb{E}_{(s,a)\sim\mathcal{D}_{\mu},(r,s^{\prime})\sim P_{\bar{\phi}}(\cdot|s,a)}\left[\log P_{\bar{\phi}}(r,s^{\prime}|s,a)-\log P_{\phi^{*}}(r,s^{\prime}|s,a)\right]=\mathbb{E}_{(s,a)\sim\mathcal{D}_{\mu}}\left[KL(P_{\hat{\phi}}(\cdot|s,a)||P_{\phi^{*}}(\cdot|s,a))\right]$. According to (Mardia et al., 2020), when $3\leq K\leq\frac{N_{sa}C_{0}}{e}+2$:
@@
Appendix D Proof of Theorem 3.3
We begin by listing the key lemmas used in the proof. The first lemma is from (Laurent and Massart, 2000), which claims:
Lemma D.1.
@@
Let $X_{1}=\frac{(\mu-\hat{\mu})^{2}}{\sigma^{2}/n}$, then $X_{1}\sim\chi^{2}_{1}$. According to Lemma D.1, $\mathbb{P}(X\geq n+2\sqrt{nt}+2t)\leq\exp(-t),\ \forall t>0$. Let $\exp(-t)=\frac{\delta^{\prime}}{2}$, we obtain:
@@
Further, we observe that $\frac{\hat{\sigma}^{2}}{\sigma^{2}}=\frac{\sum_{i=1}^{n}(x_{i}-\hat{\mu})^{2}}{n\sigma^{2}}=\frac{X_{2}}{n}$, where $X_{2}\sim\chi^{2}_{n-1}$. Applying the two formulas in Lemma D.1 and letting $\exp(-t)=\delta^{\prime}/4$, we have:
@@
Since $P_{\bar{\phi}}$ is an MLE of $P_{\phi^{*}}$, $\bar{\mu}_{sa}^{i}=\hat{\mu}_{sa}^{i}$ and $\bar{\sigma}_{sa}^{2,i}=\hat{\sigma}_{sa}^{2,i}$ are the sample mean and sample variance, respectively. Applying Lemma D.2 and letting $\delta^{\prime}=\frac{\delta}{2\mathcal{D}_{sa}^{\mu}(d+1)}$, we obtain: